OpenStudy (anonymous):

z= (1+7i)/(2-i)^2 then amp(z)=?

OpenStudy (shamim):

|2-i|^2=?

OpenStudy (unklerhaukus):

\[z= \frac{1+7i}{(2-i)^2} \] start by expanding the square in the denominator \[(2-i)^2=(2-i)\times(2-i)=\]

OpenStudy (anonymous):

then

OpenStudy (unklerhaukus):

what did you get?

OpenStudy (anonymous):

pls explain in details

OpenStudy (unklerhaukus):

\[(2−i)×(2−i)=2(2−i)-i(2−i)=\]

OpenStudy (unklerhaukus):

\[2(2-i)=2\times2+2\times-i=\]

OpenStudy (unklerhaukus):

\[i(2−i)=i\times2+i\times-i=\] [ don't forget that i × i = i^2 = -1 ]

OpenStudy (anonymous):

then

OpenStudy (unklerhaukus):

show me your working

OpenStudy (anonymous):

i can't understand pls solve it in details

OpenStudy (unklerhaukus):

booo

OpenStudy (dumbcow):

if you just want solution ... use wolfram http://www.wolframalpha.com/input/?i=%281%2B7i%29%2F%282-i%29%5E2

OpenStudy (unklerhaukus):

boooo

OpenStudy (unklerhaukus):

\[z\\= \frac{1+7i}{(2-i)^2}\\=\frac{1+7i}{(2−i)(2−i)}\\=\frac{1+7i}{2(2−i)-i(2−i)} \\=\frac{1+7i}{2×2+2×−i-(i×2+i×−i)}\\=\frac{1+7i}{4-2i-(2i-i^2)}\\=\color{white}{\frac{1+7i}{4-2i-2i-1}}\\\color{white}{=\frac{1+7i}{3-4i}}\] after you have simplified this denominator completely use the complex conjugate to realize the denominator